OpenStudy (anonymous):

determine if (-5/2, 1) is a solution to -2x+3y=2 please show work of how you get the solution

OpenStudy (anonymous):

All you have to do is plug in x=5/2 and y=1 Therefore you have... -2(-5/2) + 3(1)= 5+3 != 2 Using != to mean not equal. Therefore the equation is not true.

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

the average price of a certain brand of television dropped from $3000 in january, to $2,400 in september. find the rate of change in the average price of the television.. show work please